What is the equation for line of pursuit when you start at (45,0) and go to (2,50)?
Answer by tjhd(7)

I'm not sure this is what you're looking for but....
45 = x1
2 = x2
0 = y1
50 = y2
.............
subtract them (x1 - x2) and (y1 - y2)
45 - 2 = 43
0 - 50 = -50
then put them like... x/y (x over y)
that is the X(so you get -43/50x.
Take one of the points (either (45,0) or (2,50) and put them into y=mx+b formula...
y= -43/50 (2) + 2
y = .28... then you put that into the formula as this... and your answer is...
y = -43/50x + .28
I'm sorry if this is not the answer you're looking for, but I tried to help :)